# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v2

@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We use Amber Wang et al. (2023), MSCOCO Lin et al. (2014) and LLaVa-QA90 Liu et al. (2024b) datasets for the generative task, and Amber Wang et al. (2023) and POPE Li et al. (2023b) datasets on the discriminative tasks.
@@
Implementation Details. The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling Graves (2013).
@@ 4.1 Results on Open-ended Generation
MLLMs Method Generative Discriminative Score $\uparrow$
cha $\downarrow$ cov $\uparrow$ hal $\downarrow$ cog $\downarrow$ acc $\uparrow$ $\uparrow$ $\uparrow$ F1 $\uparrow$
@@
RLAIF-V 51 6.6 49.7 32.0 2.9 76.7 92.0 78.1 84.5 89.0
VCD 26 6.7 46.4 32.6 3.5 71.3 91.1 62.3 74.3 83.8
M3ID 12 6.2 50.5 29.3 2.8 72.4 91.8 64.1 75.5 84.7
AVISC 48 6.5 50.2 34.8 2.7 73.8 89.7 68.4 77.6 85.5
@@
RLAIF-V 51 7.6 47.7 29.9 2.8 76.5 84.5 79.0 81.7 87.1
VCD 26 7.9 49.7 36.7 3.7 75.9 83.5 79.3 81.3 86.7
M3ID 12 7.3 49.2 33.8 3.7 75.8 84.4 77.9 81.0 86.9
AVISC 48 7.1 48.8 34.4 4.3 75.9 83.4 79.5 81.4 87.2
@@
RLAIF-V 51 7.8 50.5 35.7 4.1 81.2 90.8 79.7 84.9 88.6
VCD 26 8.0 51.3 35.3 4.1 75.6 83.5 78.8 81.1 86.6
M3ID 12 7.8 51.7 34.9 4.1 75.9 83.5 79.3 81.3 86.8
AVISC 48 10.9 50.5 35.5 4.4 82.1 90.7 81.4 85.8 87.5
@@
Table 2: Results Using the CHAIR Metric on the MSCOCO Subset Following Deng et al. (2024).
Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2
$C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$
base 48.8 13.4 78.6 99.8 57.8 16.5 73.6 101.3 59.2 17.6 75.8 105.3
VCD 44.8 12.8 76.8 89.8 63.4 19.6 71.2 95.5 52.7 16.1 73.2 93.6
M3ID 44.5 12.1 77.0 85.1 57.3 16.1 72.5 100.1 52.4 15.8 72.7 92.6
AVISC 46.4 13.4 76.3 90.5 58.9 17.8 70.6 99.6 58.3 17.5 75.6 99.5
Ours 41.0 10.9 77.3 85.0 47.4 13.4 72.3 93.9 45.0 12.4 74.9 83.5
Hallucination Mitigation. Our approach reduces hallucination by up to $33\%$ at sentence-level (halhal in Tab. 1 and $C_{S}$ in Tab. 2) and $32\%$ at instance-level (chacha in Tab. 1 and $C_{I}$ in Tab. 2), demonstrating its effectiveness in mitigating overall hallucinations. It also effectively mitigates co-occurrence hallucinations, with reductions of up to $57\%$ for cogcog in Tab. 1. Accuracy gains of up to $92\%$ (Tab. 3) further demonstrate that our model improves alignment with the input image, highlighting its ability to jointly address text–visual and co-occurrence biases (Sec. 4.4).
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following Leng et al. (2024).
@@ 4.2 Results on Discriminative Task
We next evaluate our method on discriminative tasks using AMBER (discriminative VQA) and POPE (existence VQA), with results shown in Tab. 1 and Tab. 4. Our approach achieves a better balance between precision and recall, yielding consistently higher F1 scores and improved accuracy. Notably, unlike competing methods that degrade Intern-VL2, ours preserves its performance via bias awareness, though improvements differ across categories and MLLMs.
@@
Method LLaVA-v1.5 InstructBLIP mPLUG-Owl2 InternVL2
$uparrow Acc\uparrow F1\uparrow$ $uparrow Acc\uparrow F1\uparrow$
@@
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using LLaVA-v1.5 Liu et al. (2024a). Our method improves performance across all categories, with the largest gains in existence, attributes, and state—categories strongly tied to visual cues, benefiting from enhanced visual token influence.
Variation in Improvement Across MLLMs. Our method achieves the most significant improvement on mPLUG-Owl2 in object existence VQA (Tab. 4) and on LLaVA-v1.5 across various VQA categories (Tab. 1 discriminative), with minor improvement observed on InternVL2. This variation in performance correlates with the baseline visual influence ratios of the MLLMs. Fig. 3(a) presents the visual influence ratios in object existence VQA, showing that LLaVA-v1.5 exhibits the lowest visual contribution, followed by mPLUG-Owl2. This lower baseline visual influence ratio allows our method to make more impactful adjustments. In contrast, InternVL2 has an original visual influence ratio exceeding $50\%$, resulting in minimal improvement when our method is applied. The strong performance of InternVL2 can be attributed to its original high visual influence ratio, further validating the motivation behind our approach. However, this dominant visual influence ratio in InternVL2 is not consistent across other question categories or in open-ended generation tasks  [^3] Results for other question categories and visualizations of open-ended generation are in supplementary., which allows our method to enhance its performance on the AMBER dataset (Tab. 6).
@@ 4.3 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6, Qwen2-VL and Intern-VL2, on the Amber dataset. As shown in Tab. 6 (supplementary), our method consistently improves performance, even for strong baselines. To test generality, we also evaluate on MMBench Liu et al. (2025), MMVet Yu et al. (2023), ScienceQA Lu et al. (2022), and VizWiz Gurari et al. (2018). Results (supplementary) show that our approach enhances perception and recognition, especially for tasks with clear visual grounding, while preserving overall task performance and demonstrating robustness.
4.4 Ablation Study
In this section, we first analyze text–visual and co‐occurrence biases, and then study the contributions of our proposed components. Further details on gradient computation methods, norm selection, and hyperparameter settings are provided in the supplementary material.
@@ 4.4 Ablation Study
Co-occurrence Bias Analysis. Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset Lin et al. (2014). Results in Fig. 5(c) indicate that when only ‘chair’ or ‘table’ exists in the image, the hallucination rate for the other object is $23.5\%$, with a $31.9\%$ rate of sharing the same most influential token. This indicates that the ‘Same Most Influential Token’ phenomenon is common in co-occurrence hallucinations. Our GACD effectively reduces the hallucination where both ‘table’ or ‘chair’ are predicted in single-object images.
Component Analysis. To assess the effectiveness of each component in our proposed method, we conducted the following variants: 1) Visual Amplification (VA) only: visual amplification is applied to all visual tokens ($\mathbf{t}^{v}$), including during noun predictions. 2) Co-occurrence Hallucination Reduction (CR): object-related visual tokens are detected, and $\mathbf{t}^{u}$ is amplified during noun predictions. 3) Our full model, with early stopping (ES). Tab. 5 demonstrates that each component of our method contributes to the overall performance. VA significantly reduces hallucinations while improving object recall. CR further mitigates co-occurrence bias, a residual form of the text-visual bias addressed by VA, resulting in additional hallucination reduction. Both VA and CR achieve these gains without introducing trade-offs. When necessary, the ES mechanism shortens outputs to effectively reduce hallucinations, with only a slight recall trade-off that remains acceptable for most MLLMs.
@@
Components LLaVA-v1.5 InstructBLIP mPLUG-Owl2
VA CR ES $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$ $C_{S}\downarrow$ $C_{I}\downarrow$ $R\uparrow$ $Len\uparrow$
48.8 13.4 78.6 99.8 57.8 16.5 73.6 101.3 59.2 17.6 75.8 105.3
46.4 11.6 79.0 95.6 53.6 15.1 75.3 108.4 52.6 14.4 78.2 95.6
46.2 11.3 79.4 95.5 53.2 14.0 74.6 105.7 52.3 14.2 78.0 95.5
41.0 10.9 77.3 85.0 47.4 13.4 72.3 93.9 45.0 12.4 74.9 83.5
@@ 5 Conclusion
Appendix A First Order Taylor Expansion
@@
Appendix B Interpreting Contrastive Decoding through KL Divergence
@@
Appendix C Result on modern MLLMs
@@ Appendix D MLLMs Architectures
Appendix E Results on MMBench
We further evaluate our method on MMBench Liu et al. (2025). The results in Tab. 8 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ Appendix E Results on MMBench
Appendix F Results on MM-Vet
The evaluation on MM-Vet Yu et al. (2023) in Tab. 9 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.

### v3

@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We use Amber [wang2023llm], MSCOCO [lin2014microsoft] and LLaVa-QA90 [liu2024visual] datasets for the generative task, and Amber [wang2023llm] and POPE [li2023evaluating] datasets on the discriminative tasks.
4.1 Evaluation Metrics
@@
4.2 Implementation Details
The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling [graves2013generating].
@@ 4.3 Results on Open-ended Generation
Method Generative Task Discriminative Task Score
cha cov hal cog acc P R F1
$\downarrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$
@@
RLAIFv yu2024rlaif 6.6 49.7 32.0 2.9 76.7 92.0 78.1 84.5 89.0
VCD leng2024mitigating 6.7 46.4 32.6 3.5 71.3 91.1 62.3 74.3 83.8
M3ID favero2024multi 6.2 50.5 29.3 2.8 72.4 91.8 64.1 75.5 84.7
AVISC woo2024don 6.5 50.2 34.8 2.7 73.8 89.7 68.4 77.6 85.5
@@
RLAIFv yu2024rlaif 7.6 47.7 29.9 2.8 76.5 84.5 79.0 81.7 87.1
VCD leng2024mitigating 7.9 49.7 36.7 3.7 75.9 83.5 79.3 81.3 86.7
M3ID favero2024multi 7.3 49.2 33.8 3.7 75.8 84.4 77.9 81.0 86.9
AVISC woo2024don 7.1 48.8 34.4 4.3 75.9 83.4 79.5 81.4 87.2
@@
RLAIFv yu2024rlaif 7.8 50.5 35.7 4.1 81.2 90.8 79.7 84.9 88.6
VCD leng2024mitigating 8.0 51.3 35.3 4.1 75.6 83.5 78.8 81.1 86.6
M3ID favero2024multi 7.8 51.7 34.9 4.1 75.9 83.5 79.3 81.3 86.8
AVISC woo2024don 10.9 50.5 35.5 4.4 82.1 90.7 81.4 85.8 87.5
@@
Table 2: Open-ended Generation Results Using the CHAIR Metric on the MSCOCO Subset Following [deng2024seeing].
Models Metrics Baseline VCD M3ID AVISC Ours
LLaVA-v1.5 $C_{S}\downarrow$ 48.8 44.8 44.5 46.4 41.0
$C_{I}\downarrow$ 13.4 12.8 12.1 13.4 10.9
$R\uparrow$ 78.6 76.8 77.0 76.3 77.3
$Len\uparrow$ 99.8 89.8 85.1 90.5 85.0
InstructBLIP $C_{S}\downarrow$ 57.8 63.4 57.3 58.9 47.4
$C_{I}\downarrow$ 16.5 19.6 16.1 17.8 13.4
$R\uparrow$ 73.6 71.2 72.5 70.6 72.3
$Len\uparrow$ 101.3 95.5 100.1 99.6 93.9
mPLUG-Owl2 $C_{S}\downarrow$ 59.2 52.7 52.4 58.3 45.0
$C_{I}\downarrow$ 17.6 16.1 15.8 17.5 12.4
$R\uparrow$ 75.8 73.2 72.7 75.6 74.9
$Len\uparrow$ 105.3 93.6 92.6 99.5 83.5
Hallucination Mitigation. Our approach reduces hallucination by up to $33\%$ at sentence-level (halhal in Tab. 1 and $C_{S}$ in Tab. 2) and $32\%$ at instance-level (chacha in Tab. 1 and $C_{I}$ in Tab. 2), demonstrating its effectiveness in mitigating overall hallucinations. It also effectively mitigates co-occurrence hallucinations, with reductions of up to $57\%$ for cogcog in Tab. 1. Accuracy gains of up to $92\%$ (Tab. 3) further demonstrate that our model improves alignment with the input image, highlighting its ability to jointly address text–visual and co-occurrence biases (Sec. 5).
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following [leng2024mitigating].
@@ 4.4 Results on Discriminative Task
We next evaluate our method on discriminative tasks using AMBER (discriminative VQA) and POPE (existence VQA), with results shown in Tab. 1 and Tab. 4. Our approach achieves a better balance between precision and recall, yielding consistently higher F1 scores and improved accuracy. Notably, unlike competing methods that degrade Intern-VL2, ours preserves its performance via bias awareness. Category-wise analysis further shows heterogeneous gains, with improvements varying across question types.
@@
Method LLaVA Instruct mPLUG Intern
v1.5 BLIP Owl2 VL2
Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ F1 $\uparrow$
@@
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using LLaVA-v1.5 [liu2024llava]. Our method improves performance across all categories, with the largest gains in existence, attributes, and state—categories strongly tied to visual cues, benefiting from enhanced visual token influence.
Variation in Improvement Across MLLMs. Our method achieves the most significant improvement on mPLUG-Owl2 in object existence VQA (Tab. 4) and on LLaVA-v1.5 across various VQA categories (Tab. 1 discriminative), with minor improvement observed on InternVL2. This variation in performance correlates with the baseline visual influence ratios of the MLLMs. Fig. 3(a) presents the visual influence ratios in object existence VQA, showing that LLaVA-v1.5 exhibits the lowest visual contribution, followed by mPLUG-Owl2. This lower baseline visual influence ratio allows our method to make more impactful adjustments. In contrast, InternVL2 has an original visual influence ratio exceeding $50\%$, resulting in minimal improvement when our method is applied. The strong performance of InternVL2 can be attributed to its original high visual influence ratio, further validating the motivation behind our approach. However, this dominant visual influence ratio in InternVL2 is not consistent across other question categories or in open-ended generation tasks  [^3] Results for other question categories and visualizations of open-ended generation are in supplementary., which allows our method to enhance its performance on the AMBER dataset (Tab. 1 in supplementary).
@@ 4.5 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6, Qwen2-VL and Intern-VL2, on the Amber dataset. As shown in Tab. 1 in supplementary, our method consistently improves performance, even for strong baselines. To further evaluate the effectiveness of our method across diverse tasks, we conduct evaluations on the MMBench [liu2025mmbench], MMVet [yu2023mm], ScienceQA [lu2022learn] and VizWiz [gurari2018vizwiz] datasets. Results in supplementary show that our approach enhances perception and recognition, especially for tasks with clear visual grounding, while preserving overall task performance and demonstrating robustness.
5 Ablation Study
@@
In this section, we first analyze text–visual and co‐occurrence biases, then evaluate the contributions of our proposed components. We also detail the gradient-computation methods and norm selection. Additional ablation studies and hyperparameter settings are provided in the Supplementary Material.
5.1 Text-Visual Bias Analysis
@@
5.2 Co-occurrence Bias Analysis
Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. These visualizations indicate that the influence distribution over tokens is typically sharply peaked, so selecting the single most influential token gives a clean and interpretable attribution signal while reducing noise from low-influence tokens. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset [lin2014microsoft].
Results in Fig. 5(c) show that when only a ‘chair’ or ‘table’ exists in the image, the other object is hallucinated in $23.5\%$ of cases, with $31.9\%$ sharing the same most influential token, indicating that the ‘Same Most Influential Token’ phenomenon is common in co-occurrence hallucinations. Our GACD effectively reduces such hallucination where both ‘table’ or ‘chair’ are predicted in single-object images.
5.3 Component Analysis
To assess the effectiveness of each component in our proposed method, we conducted the following variants: 1) Visual Amplification (VA) only: visual amplification is applied to all visual tokens ($\mathbf{t}^{v}$), including during noun predictions. 2) Co-occurrence Hallucination Reduction (CR): object-related visual tokens are detected, and $\mathbf{t}^{u}$ is amplified during noun predictions. 3) Our full model, with early stopping (ES). Tab. 5 demonstrates that each component of our method contributes to the overall performance. VA significantly reduces hallucinations while improving object recall. CR further mitigates co-occurrence bias, a residual form of the text-visual bias addressed by VA, resulting in additional hallucination reduction. Both VA and CR achieve these gains without introducing trade-offs. When necessary, the ES mechanism shortens outputs to effectively reduce hallucinations, with only a slight recall trade-off.
@@ 5.3 Component Analysis
Models VA
CR
ES
LLaVA-v1.5 $C_{S}\downarrow$ 48.8 46.4 46.2 41.0
$C_{I}\downarrow$ 13.4 11.6 11.3 10.9
$R\uparrow$ 78.6 79.0 79.4 77.3
$Len\uparrow$ 99.8 95.6 95.5 85.0
InstructBLIP $C_{S}\downarrow$ 57.8 53.6 53.2 47.4
$C_{I}\downarrow$ 16.5 15.1 14.0 13.4
$R\uparrow$ 73.6 75.3 74.6 72.3
$Len\uparrow$ 101.3 108.4 105.7 93.9
mPLUG-Owl2 $C_{S}\downarrow$ 59.2 52.6 52.3 45.0
$C_{I}\downarrow$ 17.6 14.4 14.2 12.4
$R\uparrow$ 75.8 78.2 78.0 74.9
$Len\uparrow$ 105.3 95.6 95.5 83.5
5.4 Gradient Computation and Norm Analysis
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and enabling straightforward reproducibility. For comparison, we evaluate Integrated Gradients (IG) [enguehard2023sequential, lundstrom2022rigorous, kapishnikov2021guided] using the SIG [enguehard2023sequential] implementation; results for this ablation on the POPE MSCOCO adversarial setting are shown in Tab. 6. ‘IG’ denotes the SIG-based results, while “Ours” refers to our direct gradient method. Both achieve comparable accuracy and F1‐score, but the direct-gradient variant is substantially more efficient.
Table 6: Gradient Methods on the POPE MSCOCO Dataset(Adv.)
Methods Accuracy F1 Speed(ms)
MPLUG-Owl2 IG [enguehard2023sequential] 83.4 82.9 20335
Ours 84.2 83.7 385
[…]
Table 7: Norm Strategies on the POPE MSCOCO Dataset(Adv.)
Norm LLaVA-v1.5 InstructBLIP mPLUG-Owl2
Acc F1 Acc F1 Acc F1
L1 83.5 82.1 82.5 82.1 84.2 83.7
L2 83.2 81.9 79.5 79.6 83.2 82.9
L 83.4 82.0 82.1 81.8 80.8 80.6
@@ 6 Conclusion
Limitations. Our method is limited to white-box MLLMs, as it requires access to gradients. Its effectiveness depends on the baseline MLLM’s original visual influence ratio, and the importance of visual information. For instance, existence questions are primarily rely on visual, whereas relational questions are less direct and require inference beyond visual. As a post-processing technique, our method does not involve model training. In future work, we aim to explore how insights from GACD can guide and improve training strategies for enhanced visual perception in MLLMs.
[…]
Supplementary Material
1 First Order Taylor Expansion
@@
2 Interpreting Contrastive Decoding through KL Divergence
@@
3 Result on modern MLLMs
@@ 4 MLLMs Architectures
5 Results on MMBench
We further evaluate our method on MMBench [liu2025mmbench]. The results in Tab. 3 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ 5 Results on MMBench
6 Results on MM-Vet
The evaluation on MM-Vet [yu2023mm] in Tab. 4 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.